OpenStudy (anonymous):

find f(2), given f(x)=4x-1

OpenStudy (08surya):

put x=2

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ f(x)=4x-1}\) For example f(3): \(\normalsize\color{black}{ f(3)=4(3)-1=12-1=11}\)

OpenStudy (solomonzelman):

Can you do the f(2) just the same way ?

OpenStudy (solomonzelman):

btw- actuality is a good word :)

OpenStudy (anonymous):

f(-2) = -4(-2) - 1 = 8 - 1 = 7